# Magnetic Feshbach resonances in ultracold atom-molecule collisions

### v1

@@ Abstract
The Rb-SrF collision complex is characterized by two deep and strongly anisotropic PESs of singlet and triplet symmetries [52], which can couple hundreds of rotational states [45], in addition to the pronounced hyperfine structure of both SrF and Rb, causing major computational challenges, as mentioned above. Here, we address these challenges using a recently developed total rotational angular momentum (TRAM) representation [53], which enables us to efficiently incorporate all rotational, fine, and hyperfine states of the collision partners in converged CC calculations on realistic ab initio PESs. Our calculated MFR spectra are dominated by broad resonances with a density of about 5 resonances per 1 kG, which should be readily observable experimentally. Our results open up the possibility of rigorous CC calculations of MFRs on a wide range of ultracold atom-molecule collisions of experimental interest.
@@
where $\mu$ is the reduced mass of the atom-molecule trimer described by the Jacobi vectors $\mathbf{R}$ (pointing from the molecule’s center of mass to the atom) and $\mathbf{r}$ (joining the nuclei of the diatomic molecule), and $\hat{L}^{2}$ is the squared orbital angular momentum for the collision. The atom-molecule interaction $\hat{V}_{\text{int}}=\sum_{S,M_{S}}V_{S}(R,\theta)|SM_{S}\rangle\langle SM_{S}|+\hat{V}_{\text{mdd}}$, where $V_{S}(R,\theta)$ are the singlet ($S=0$) and triplet ($S=1$) adiabatic PESs (see below), and $\hat{V}_{\text{mdd}}$ and the magnetic dipole-dipole interaction. We adopt the rigid-rotor approximation by setting $r=r_{e}$, the equilibrium distance in SrF, which is expected to be qualitatively accurate for collisions of rigid molecules such as SrF [55]. As such, the atom-molecule PES only depends on $R$ and $\theta$, the angle between $\mathbf{R}$ and $\mathbf{r}$. In Eq. 1, $\hat{H}_{\text{mol}}$ describes the isolated ${}^{2}\Sigma$ molecule, $\hat{H}_{\text{mol}}=\hat{H}_{\text{mol}}^{\text{rot}}+\hat{H}_{\text{mol}}^{\text{spin}}+\hat{H}_{\text{mol}}^{\text{spin-rot}}$, where $\hat{H}_{\text{mol}}^{\text{rot}}=B_{e}\hat{\mathbf{N}}^{2}$ is the rotational Hamiltonian, $B_{e}$ is the rotational constant, $\hat{\mathbf{N}}$ is the rotational angular momentum of the diatomic molecule, and $\hat{H}_{\text{mol}}^{\text{spin}}=g_{S}\mu_{0}B\hat{S}_{Z}+a\hat{\mathbf{I}}\cdot\hat{\mathbf{S}}$ is the spin Hamiltonian, which accounts for the electron and nuclear spins within the molecule described by the operators $\hat{\mathbf{S}}$ and $\hat{\mathbf{I}}$ interacting with each other via the Fermi contact interaction (coupling constant $a$) and with an external magnetic field $B$ directed along a space-fixed $z$-axis. The spin-rotation interaction $\hat{H}_{\text{mol}}^{\text{spin-rot}}=\gamma_{\text{sr}}\hat{\mathbf{N}}\cdot\hat{\mathbf{S}}+\frac{c\sqrt{6}}{3}\bigl{(}\frac{4\pi}{5}\bigr{)}^{1/2}\sum_{q}(-1)^{q}Y_{2-q}(\theta_{r},\phi_{r})[\hat{\mathbf{I}}\otimes\hat{\mathbf{S}}]^{(2)}_{q}$ includes the electron spin-rotation and anisotropic hyperfine interactions, parametrized by the coupling constants $\gamma_{\text{sr}}$ and $c$. The internal structure of atomic Rb(2S) is described by the Hamiltonian $\hat{H}_{\text{at}}=g_{S}\mu_{0}B\hat{S_{a}}_{z}+A_{a}\hat{\mathbf{I}}_{a}\cdot\hat{\mathbf{S}}_{a}$, where ${\mathbf{S}}_{a}$ and $\hat{\mathbf{I}}_{a}$ are the atomic electron and nuclear spin operators, and $A_{a}$ is the hyperfine constant.
@@
To map out the spectrum of MFRs in ultracold Rb + SrF collisions, we solve the Schrödinger equation $\hat{H}|\Psi\rangle=E|\Psi\rangle$, where $E$ is the total energy and $|\Psi\rangle$ is the scattering wavefunction, by expanding the wavefunction of the atom-molecule collision complex as $|\Psi\rangle=R^{-1}\sum_{i}F_{i}(R)|\Phi_{i}\rangle$. We use a recently developed total rotational angular momentum (TRAM) basis ${|\Phi_{i}\rangle}=|(Nl)J_{r}M_{r}\rangle|n_{s}\rangle|n_{s}^{(a)}\rangle$, which enables a computationally efficient treatment of the essential hyperfine structure of the collision partners interacting via highly anisotropic potentials in an external magnetic field [53]. Here, $|(Nl)J_{r}M_{r}\rangle$ are the eigenstates of $\hat{J}_{r}^{2}$ and $\hat{J}_{z}$, where $\hat{\mathbf{J}}_{r}=\mathbf{N}+\mathbf{L}$ is the TRAM of the collision complex and $\hat{J}_{z}$ is its projection on the field axis. The atomic and molecular spin functions are denoted by $|n_{s}^{(a)}\rangle$ and $|n_{s}\rangle$.
@@
Unlike the total angular momentum of the collision pair, $J_{r}$ is rigorously conserved even in the presence of external magnetic fields and isotropic hyperfine interactions [53]. The different values of $J_{r}$ can only be coupled by weak spin-rotation interactions, such as the spin-rotation and anisotropic hyperfine interactions [53]. The ensuing near-conservation of $J_{r}$ is key to the computational efficiency of the TRAM basis, which enables the inclusion of a sufficient number of closed rotational channels ($N_{\text{max}}=175$ [56]) to produce converged results for highly anisotropic Rb + SrF collisions. We integrate the CC equations numerically using the log-derivative method [57, 58] to obtain the asymptotic behavior of the radial solutions $F_{i}(R)$, from which the $S$-matrix elements and scattering cross sections are obtained as a function of collision energy and magnetic field as described in Ref. [53] (see also the Supplemental Material [56]).
@@
Ab initio calculations. To describe the quantum dynamics of ultracold Rb + SrF collisions it is essential to have a realistic description of both the singlet and triplet PESs. Here, we employ the state-of-the-art ab initio triplet PES [45] computed using the coupled cluster method with a large basis set. To determine the singlet PES, we calculated the singlet-triplet energy splitting of Rb-SrF using the Complete Active Space Self-Consistent Field (CASSCF) method using an aug-cc-pVQZ basis set and ECP28MDF core potentials for Rb and Sr. The active space for this calculation includes the open-shell orbitals of both Rb and SrF, accommodating 2 electrons in 2 orbitals. The exchange energy thus obtained is added to the triplet PES [45] to produce the singlet PES. This procedure ensures the correct behavior of the singlet PES in the limit $R\to\infty$, where the exchange energy decays exponentially, and accurately accounts for dispersion interactions, which should asymptotically be identical for both the singlet and triplet PESs. We verified the correctness of the singlet-triplet energy splitting using additional spin-flip Symmetry Adapted Perturbation Theory (SAPT) calculations [59], in which the asymptotic behavior is calculated based on monomer properties [56].
@@
Magnetic Feshbach resonances in ultracold atom-molecule collisions. We consider ultracold Rb atoms and SrF molecules colliding in the initial states |$\ket{i}\otimes\ket{j}$, where |$\ket{i}$ and |$\ket{j}$ stand for the initial hyperfine-Zeeman states of Rb and SrF, labeled in the order of increasing energy, as shown in Fig. 1(c). In the low $B$ field limit, these states are identical to the atomic and molecular hyperfine states |$\ket{F_{a}m_{F_{a}}}$ and |$\ket{(NSI)Fm_{F}}$, where $\hat{\mathbf{F}}=\hat{\mathbf{N}}+\hat{\mathbf{S}}+\hat{\mathbf{I}}$ is the total angular momentum of SrF, which is a vector sum of the rotational, electron spin, and nuclear spin angular momenta. We are interested in the experimentally relevant case of SrF molecules colliding in their ground rotational states $(N=0)$, which can be prepared by laser cooling and trapped in an optical dipole trap [32, 35]. The hyperfine structure of SrF($N=0$) consists of two hyperfine manifolds with $F=1,0$, which are split by an external magnetic field into 4 levels [see Fig. 1(c)]. Non-fully spin-polarized initial states of Rb and SrF can undergo collision-induced spin-exchange, i.e., |$\ket{6}\otimes\ket{2}$to |$\ket{3}\otimes\ket{3}$, which conserves the projection of the total internal angular momentum of the collision partners $M_{F}=m_{F_{1}}+m_{F_{2}}$. The rigorous treatment of the spin-exchange collisions and of the hyperfine structure of the collision partners constitute two key improvements over the previous calculations [45].
@@
In Fig. 2(a), we show the calculated spectrum of MFRs in ultracold Rb + SrF collisions obtained from fully converged CC calculations with 1232 TRAM basis functions. The resonance density (4-5 per 1 kG) is fairly insensitive to the initial and final hyperfine states of Rb and SrF, and is comparable to that of MFRs in ultracold Rb + Rb collisions [60] The hyperfine structure of the collision partners and their short-range spin-exchange interactions play a crucial role in inducing the MFRs: When the spin-exchange coupling is omitted in test calculations, most of the MFRs shown in Fig. 2(a) disappear [56].
@@
Figure 2: (a) Calculated MFR spectra for elastic and inelastic Rb($|6\rangle$) + SrF($|2\rangle$) collisions at $E_{\text{c}}=0.1\,\mu$K. The elastic cross section (black dashed line) is multiplied by 1000 for visual clarity. The final states are marked as ($i,j$), where $i$ and $j$ label the internal states of Rb and SrF (see text and Fig. 1(c)). (b) Same as (a) but with only the $N=0$ rotational of SrF retained in the CC basis.
To further elucidate the role of anisotropic couplings between different rotational states, we diagonalized the atom-molecule Hamiltonian in Eq. 1 without the radial kinetic energy term to produce the adiabatic states of the Rb-SrF collision complex as a function of $R$. In the adiabatic picture, the Rb + SrF collision partners approach each other at long range on a smooth adiabatic potential, which experiences a series of avoided crossings with other potentials as $R$ is decreased, leading to inelastic transitions and MFRs [56]. The density of the adiabatic potentials in the vicinity of and below the incident collision threshold is fairly well-converged already for moderate values of $N_{\text{max}}\simeq 60$. Further increasing $N_{\text{max}}$ only produces slight changes in the shapes of the existing adiabatic curves, leading to small changes in closed-channel bound states, which are not enough to produce new bound states that could result in additional MFRs. Therefore, the MFRs density saturates at $N_{\text{max}}\simeq 60$, well below the value required for full convergence ($N_{\text{max}}=175$). In other words, even though the adiabatic states themselves are not completely converged at low $N_{\text{max}}$, their density near the incident threshold is. This explains why highly excited rotational states do not contribute directly to the MFR density, although they do affect the positions and widths of the individual MFRs.
@@
It has been hypothesized that MFRs in ultracold collisions of alkali-dimers with alkali-metal atoms are caused by the long-range LeRoy-Bernstein bound states [61], which are only weakly affected by the atom-molecule interaction anisotropy, since it typically decreases much faster with $R$ than the isotropic part of the PES. If this hypothesis is true, the atom-molecule MFRs would be very similar to the well-known atom-atom MFRs [2] in that the effects of short-range interaction anisotropy can be safely neglected. An opposing hypothesis states that MFS are due to the short-range chaotic states of the atom-molecule collision complex [62, 63]
To explore the short-range vs. long-range character of atom-molecule MFRs, we performed additional CC calculations with the short-range part of the Rb-SrF PESs scaled by a factor of $\lambda=0.5$ (see [56] for details). Because the long-range parts of the scaled and unscaled PESs are the same, the scaling only changes the density of short-range complex states. As a result, it should not modify the density of MFRs if, as proposed in Ref. [61], the MFRs are due to the long-range states. If, on the other hand, the MFRs are due to the short-range states, then the PES scaling is expected to lead to a substantial decrease in MFR density according to the approximate scaling $\rho\simeq(\lambda)^{3/2}$ [64]. As shown in Fig. 3, the number of MFRs decreases by 40% upon shallowing the PESs, suggesting that MFRs in Rb + SrF collisions are predominantly caused by short-range complex states. We emphasize that these results do not completely rule out the possibility that some MFRs could be caused by long-range states, which can have much narrower widths than those caused by the short-range complex states [56], warranting further calculations and experimental measurements on near-threshold bound states of Rb-SrF, whose zero-energy crossings with collision thresholds give rise to MFRs [2]. Such calculations could be enabled by combining the numerically efficient TRAM basis used here with established techniques for solving the bound-state Schrödinger equation [65].

### v2

@@ Abstract
The Rb-SrF collision complex is characterized by two deep and strongly anisotropic potential energy surfaces (PESs) of singlet and triplet symmetries [52], which can couple hundreds of rotational states [45], in addition to the pronounced hyperfine structure of both SrF and Rb, causing major computational challenges, as mentioned above. Here, we address these challenges using a recently developed total rotational angular momentum (TRAM) representation [53], which enables us to efficiently incorporate all rotational, fine, and hyperfine states of the collision partners in converged CC calculations on realistic ab initio PESs. Our calculated MFR spectra are dominated by broad resonances with a density of about 5 resonances per 1 kG, which should be readily observable experimentally. Our results open up the possibility of rigorous CC calculations of MFRs on a wide range of ultracold atom-molecule collisions of experimental interest.
@@
where $\mu$ is the reduced mass of the atom-molecule trimer described by the Jacobi vectors $\mathbf{R}$ (pointing from the molecule’s center of mass to the atom) and $\mathbf{r}$ (joining the nuclei of the diatomic molecule), and $\hat{L}^{2}$ is the squared orbital angular momentum for the collision. The atom-molecule interaction is given as $\hat{V}_{\text{int}}=\sum_{S,M_{S}}V_{S}(R,\theta)|SM_{S}\rangle\langle SM_{S}|+\hat{V}_{\text{mdd}}$, where $V_{S}(R,\theta)$ are the singlet ($S=0$) and triplet ($S=1$) adiabatic PESs (see below), and $\hat{V}_{\text{mdd}}$ is the magnetic dipole-dipole interaction. We adopt the rigid-rotor approximation by setting $r=r_{e}$, the equilibrium distance in SrF, which is expected to be qualitatively accurate for collisions of rigid molecules such as SrF [55]. As such, the atom-molecule PES only depends on $R$ and $\theta$, the angle between $\mathbf{R}$ and $\mathbf{r}$. In Eq. 1, $\hat{H}_{\text{mol}}$ describes the isolated ${}^{2}\Sigma$ molecule, $\hat{H}_{\text{mol}}=\hat{H}_{\text{mol}}^{\text{rot}}+\hat{H}_{\text{mol}}^{\text{spin}}+\hat{H}_{\text{mol}}^{\text{spin-rot}}$, where $\hat{H}_{\text{mol}}^{\text{rot}}=B_{e}\hat{\mathbf{N}}^{2}$ is the rotational Hamiltonian, $B_{e}$ is the rotational constant, $\hat{\mathbf{N}}$ is the rotational angular momentum of the diatomic molecule, and $\hat{H}_{\text{mol}}^{\text{spin}}=g_{S}\mu_{0}B\hat{S}_{Z}+a\hat{\mathbf{I}}\cdot\hat{\mathbf{S}}$ is the spin Hamiltonian, which accounts for the electron and nuclear spins within the molecule described by the operators $\hat{\mathbf{S}}$ and $\hat{\mathbf{I}}$ interacting with each other via the Fermi contact interaction (coupling constant $a$) and with an external magnetic field $B$ directed along a space-fixed $z$-axis. The spin-rotation interaction $\hat{H}_{\text{mol}}^{\text{spin-rot}}=\gamma_{\text{sr}}\hat{\mathbf{N}}\cdot\hat{\mathbf{S}}+\frac{c\sqrt{6}}{3}\bigl{(}\frac{4\pi}{5}\bigr{)}^{1/2}\sum_{q}(-1)^{q}Y_{2-q}(\theta_{r},\phi_{r})[\hat{\mathbf{I}}\otimes\hat{\mathbf{S}}]^{(2)}_{q}$ includes the electron spin-rotation and anisotropic hyperfine interactions, parametrized by the coupling constants $\gamma_{\text{sr}}$ and $c$. The internal structure of atomic Rb(2S) is described by the Hamiltonian $\hat{H}_{\text{at}}=g_{S}\mu_{0}B\hat{S_{a}}_{z}+A_{a}\hat{\mathbf{I}}_{a}\cdot\hat{\mathbf{S}}_{a}$, where ${\mathbf{S}}_{a}$ and $\hat{\mathbf{I}}_{a}$ are the atomic electron and nuclear spin operators, and $A_{a}$ is the hyperfine constant.
@@
To map out the spectrum of MFRs in ultracold Rb + SrF collisions, we solve the Schrödinger equation $\hat{H}|\Psi\rangle=E|\Psi\rangle$, where $E$ is the total energy and $|\Psi\rangle$ is the scattering wavefunction, by expanding the wavefunction of the atom-molecule collision complex as $|\Psi\rangle=R^{-1}\sum_{i}F_{i}(R)|\Phi_{i}\rangle$, where ${|\Phi_{i}\rangle}=|(Nl)J_{r}M_{r}\rangle|n_{s}\rangle|n_{s}^{(a)}\rangle$ are the TRAM basis functions [56], which enable a computationally efficient treatment of the essential hyperfine structure of the collision partners interacting via highly anisotropic potentials in an external magnetic field [53]. Here, $|(Nl)J_{r}M_{r}\rangle$ are the eigenstates of $\hat{J}_{r}^{2}$ and $\hat{J}_{z}$, where $\hat{\mathbf{J}}_{r}=\mathbf{N}+\mathbf{L}$ is the TRAM of the collision complex and $\hat{J}_{z}$ is its projection on the field axis. The atomic and molecular spin functions are denoted by $|n_{s}^{(a)}\rangle$ and $|n_{s}\rangle$.
@@
Unlike the total angular momentum of the collision pair, $J_{r}$ is rigorously conserved even in the presence of external magnetic fields and isotropic hyperfine interactions [53]. The different values of $J_{r}$ can only be coupled by weak spin-rotation interactions, such as the spin-rotation and anisotropic hyperfine interactions [53]. The ensuing near-conservation of $J_{r}$ is key to the computational efficiency of the TRAM basis, which enables the inclusion of a sufficient number of closed rotational channels ($N_{\text{max}}=175$ [57]) to produce converged results for highly anisotropic Rb + SrF collisions. We integrate the CC equations numerically [58, 59] to obtain the asymptotic behavior of the radial solutions $F_{i}(R)$, from which the $S$-matrix elements and scattering cross sections are obtained as a function of magnetic field [57, 53].
@@
Ab initio calculations. To describe the quantum dynamics of ultracold Rb + SrF collisions it is essential to have a realistic description of both the singlet and triplet PESs. Here, we employ the state-of-the-art ab initio triplet PES [45] computed using the coupled cluster method with a large basis set. To determine the singlet PES, we employ the Complete Active Space Self-Consistent Field (CASSCF) method using an aug-cc-pVQZ basis set and ECP28MDF core potentials for Rb and Sr [57]. The active space for this calculation includes the open-shell orbitals of both Rb and SrF, accommodating two electrons in two orbitals. The exchange energy thus obtained is added to the triplet PES [45] to produce the singlet PES. This procedure ensures the correct behavior of the singlet PES in the limit $R\to\infty$, where the exchange energy decays exponentially, and accurately accounts for dispersion interactions, which should asymptotically be identical for both the singlet and triplet PESs. We verified the correctness of the singlet-triplet energy splitting using additional spin-flip Symmetry Adapted Perturbation Theory (SAPT) calculations [60], in which the asymptotic behavior is calculated based on monomer properties [57].
@@
Magnetic Feshbach resonances in ultracold atom-molecule collisions. We consider ultracold Rb atoms and SrF molecules colliding in the initial states |$\ket{i}\otimes\ket{j}$, where |$\ket{i}$ and |$\ket{j}$ stand for the initial hyperfine-Zeeman states of Rb and SrF, labeled in the order of increasing energy, as shown in Fig. 1(c). In the low $B$ field limit, these states are identical to the atomic and molecular hyperfine states |$\ket{F_{a}m_{F_{a}}}$ and |$\ket{(NSI)Fm_{F}}$, where $\hat{\mathbf{F}}=\hat{\mathbf{N}}+\hat{\mathbf{S}}+\hat{\mathbf{I}}$ is the total angular momentum of SrF, which is a vector sum of the rotational, electron spin, and nuclear spin angular momenta. We are interested in the experimentally relevant case of SrF molecules colliding in their ground rotational states $(N=0)$, which can be prepared by laser cooling and trapped in an optical dipole trap [32, 35]. The hyperfine structure of SrF($N=0$) consists of two hyperfine manifolds with $F=1,0$, which are split by an external magnetic field into 4 levels [see Fig. 1(c)]. Non-fully spin-polarized initial states of Rb and SrF can undergo collision-induced spin-exchange, i.e., |$\ket{6}\otimes\ket{2}$to |$\ket{3}\otimes\ket{3}$, which conserves the projection of the total internal angular momentum of the collision partners $m_{F_{a}}+m_{F}$. The rigorous treatment of the spin-exchange collisions and of the hyperfine structure of the collision partners constitute two key improvements over the previous calculations [45].
@@
In Fig. 2(a), we show the calculated spectrum of MFRs in ultracold Rb + SrF collisions obtained from fully converged CC calculations with 1232 TRAM basis functions. The inelastic and elastic cross sections are comparable, which indicates that significant inelastic losses can occur near MFRs, making it hard to prepare a stable atom-molecule mixture. Further work is needed to identify the parameter regimes, where these losses are suppressed. The resonance density (4-5 per 1 kG) is fairly insensitive to the initial and final hyperfine states of Rb and SrF, and is comparable to that of MFRs in ultracold Rb + Rb collisions [61]. The $J_{r}$-conserving MFRs are broad and similar to those which occur in ultracold collisions of alkali-metal atoms [2]. This is consistent with a recent theoretical prediction [62] of alkali-metal-like MFRs in ultracold Rb + CaF collisions based on model CC calculations, which neglected the atom-molecule interaction anisotropy. However, as shown below, the anisotropy has a non-negligible effect on MFR spectra. The hyperfine structure of the collision partners and their spin-exchange interactions also play a crucial role: When the spin-exchange coupling is omitted in test calculations, most of the MFRs shown in Fig. 2(a) disappear [57].
@@
Figure 2: (a) Calculated MFR spectra for elastic and inelastic Rb($|6\rangle$) + SrF($|2\rangle$) collisions at $E_{\text{c}}=0.1\,\mu$K. The elastic cross section (black dashed line) is multiplied by 1000 for visual clarity. The final states are marked as ($i,j$), where $i$ and $j$ label the internal states of Rb and SrF (see text and Fig. 1(c)). (b) Same as (a) but with only the $N=0$ rotational state of SrF retained in the CC basis.
To further elucidate the role of anisotropic couplings between different rotational states, we diagonalized the atom-molecule Hamiltonian in Eq. 1 without the radial kinetic energy term to produce the adiabatic states of the Rb-SrF collision complex as a function of $R$. In the adiabatic picture, the Rb + SrF collision partners approach each other at long range on a smooth adiabatic potential, which experiences a series of avoided crossings with other potentials as $R$ is decreased, leading to inelastic transitions and MFRs [57]. The density of the adiabatic potentials in the vicinity of the incident collision threshold is fairly well-converged already for moderate values of $N_{\text{max}}\simeq 30{-}60$. Further increasing $N_{\text{max}}$ only produces slight changes in the shapes of the existing adiabatic curves, leading to small changes in closed-channel bound states, which are not enough to produce new bound states that could result in additional MFRs. Therefore, the MFR density saturates at $N_{\text{max}}\simeq 30{-}60$ [57], well below the value required for full convergence ($N_{\text{max}}=175$). In other words, even though the adiabatic states themselves are not completely converged at low $N_{\text{max}}$, their density near the incident threshold is. This explains why highly excited rotational states do not contribute directly to the MFR density, although they do affect the positions and widths of the individual MFRs.
@@
It has been hypothesized that MFRs in ultracold collisions of alkali-dimers with alkali-metal atoms are caused by the long-range LeRoy-Bernstein bound states [63], which are only weakly affected by the atom-molecule interaction anisotropy, since it typically decreases much faster with $R$ than the isotropic part of the PES. If this hypothesis is true, the atom-molecule MFRs would be very similar to the well-known atom-atom MFRs [2] in that the effects of short-range interaction anisotropy can be safely neglected. An opposing hypothesis states that MFRs are due to the short-range chaotic states of the atom-molecule collision complex [64, 65].
We emphasize that all calculations on atom-molecule MFRs reported thus far (see, e.g., [62, 24]) were performed with severely restricted rotational basis sets. They were thus affected by uncontrolled basis set truncation errors, which are absent in our calculations. For example, in the calculations of Ref. [62], only the ground rotational state was included in the basis, which amounts to neglecting the rotational structure of the molecule and the crucial anisotropy of the atom-molecule interaction at short range. Because of these drastic approximations, the previous calculations have only been able to formulate qualitative hypotheses (but not quantitative predictions) regarding the properties of atom-molecule MFRs. By contrast, our calculations treat the atom-molecule interaction anisotropy and the hyperfine structure of the collision partners exactly using converged rotational basis sets. This not only enables testing the hypotheses formulated previously, but also makes it possible to predict the properties of MFRs in atom-molecule systems with realistic interactions. Our calculations also predict a new flavor of MFR due to the intramolecular spin-rotation interaction [see Fig. 1].
To explore the short-range vs. long-range character of atom-molecule MFRs, we performed additional CC calculations with the short-range part of the Rb-SrF PESs scaled by a factor of $\lambda=0.5$ (see [57] for details). Because the long-range parts of the scaled and unscaled PESs are the same, the scaling only changes the density of short-range complex states. As a result, it should not modify the density of MFRs if, as proposed in Ref. [63], the MFRs are due to the long-range states. If, on the other hand, the MFRs are due to the short-range states, then the PES scaling is expected to lead to a substantial decrease in MFR density according to the approximate scaling $\rho\simeq(\lambda)^{3/2}$ [66]. As shown in Fig. 3, the number of MFRs decreases by 40% upon shallowing the PESs, suggesting that MFRs in Rb + SrF collisions are predominantly caused by short-range complex states. This observation does not rule out the possibility that some MFRs could be caused by long-range states, which can have much narrower widths than those caused by the short-range complex states [57], warranting further calculations and experimental measurements on near-threshold bound states of Rb-SrF, whose zero-energy crossings with collision thresholds give rise to MFRs [2]. Such calculations could be enabled by combining the efficient TRAM basis used here [53] with established techniques for solving the bound-state Schrödinger equation [67].